A right triangle has one leg measuring "x" inches and second leg measuring (2x+2)inches. The hypotenuse measures (2x+3) inches long. Find the length of all three sides.
Answer by mananth(16949)

using the pythagoras theorem
x^2+(2x+2)^2=(2x+3)^2
x^2+4x^2+8x+4=4x^2+12x+9
x^2-4x-5=0
x^2-5x+x-5=0
x(x-5)+1(x-5)=0
(x-5)(x+1)=0
x= 5
The sides are
one leg 5 in
second (2x+2)===>12 in
hypotenuse (2x+3)===>13 in